Find the power of engine that can fire water jet of density 1000 kg/m3
at the speed of 30m/s upward from the pipe of radius 1cm?
1
Expert's answer
2021-08-02T08:40:23-0400

Explanations & Calculations


  • Those given conditions are valid only at the pump outlet where the water is ejected at the speed of "\small 30\,ms^{-1}" and beyond that these calculations may not seem to be effective.
  • The power, as it is defined is = "\Large \frac{\Delta\text{work}}{\Delta\text{time}}"
  • And the work done at the outlet is "\small W= \small \frac{1}{2}mv^2" and lets assume that this work is done at a small duration of "\small t(s)". Then the power needed is

"\qquad\qquad
\begin{aligned}
\small P&=\small \frac{W}{t}=\frac{1}{2}\Big(\frac{m}{t}\Big)v^2
\end{aligned}"

  • That term within the parentheses is the rate of mass flow which can be deduced to be "\small Av\rho" where A is the are of the pump.
  • Then,

"\qquad\qquad
\begin{aligned}
\small P&=\small \frac{1}{2}(Av\rho)v^2\\
&=\small \frac{1}{2}Av^3\rho\\
&=\small \frac{1}{2}(\pi(0.01m)^2\times(30ms^{-1})^3\times1000kgm^{-3})\\
&=\small \bold{4241.15\,W\approx4.2kW}
\end{aligned}"